OpenStudy (anonymous):

How do I simplify the fraction (1) / (1+(x/y))? I know the answer is y/x+y but how do I get that?

OpenStudy (anonymous):

hmm... \[\huge \frac{1}{1+\frac{x}{y}}\] This?

OpenStudy (anonymous):

yea srry had to step away

OpenStudy (anonymous):

simple first take reciprocal \[\frac{ 1 }{ \frac{ x+y }{ y } }\] then multiply the denomenator of the denomenator to the numereator  then you get the answer y/(x+y)